# Breaking the Curse of Dimensionality with Distributed Neural Computation

### v1

@@ Appendix A Overview of Technical Derivation
We now overview the proof of our main result and its full technical formulation. These objectives require us to recall definitions from the analysis of metric spaces, which we do now.
A.1 Preliminary Definitions
The metric ball in $(\mathcal{X},d)$ of radius $r>0$ at $x\in\mathcal{X}$ is denoted by $\operatorname{Ball}_{(\mathcal{X},d)}(x,r)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{z\in\mathcal{X}\colon d(x,z)<r\}$.
@@ A.1 Preliminary Definitions
The metric space $(\mathcal{X},d)$ is called doubling, if there is $C\in\mathbb{N}_{+}$ for which every metric ball in $(\mathcal{X},d)$ can be covered by at most $C$ metric balls of half its radius. The smallest such constant is called $(\mathcal{X},d)$’s doubling number, and is here denoted by $C_{(\mathcal{X},d)}$. Though this definition may seem abstract at first, (Heinonen, 2001, Theorem 12.1), provides an almost familiar characterization of all doubling metric spaces; indeed, $\mathcal{K}$ is a doubling metric space if and only if it can be identified via a suitable invertible map [^1] So called quasi-symmetric maps, see (Heinonen, 2001, page 78). with a subset of some Euclidean space. Every subset of $\mathbb{R}^{n}$, for any $n\in\mathbb{N}_{+}$, is a doubling metric space; see (Robinson, 2011, Chapter 9) for details.
@@
Fix a dimension $n\in\mathbb{N}_{+}$. The doubling number of any subset of Euclidean space is [^2] See (Robinson, 2011, Lemma 9.2) and the brief computations made in the proof of (Robinson, 2011, Lemma 9.4). at most $2^{n+1}$.
In what follows, all logarithms will be taken base $2$, unless explicitly stated otherwise, i.e. $\log_{v}$ is base $v$ for a given $v\in\mathbb{N}_{+}$ and $\log\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log_{2}$. As on (Petrova & Wojtaszczyk, 2023, page 762), the radius of a subset $A\subseteq\mathbb{R}^{n}$, denoted by $\operatorname{rad}(A)$, is defined by
$\operatorname{rad}(A)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\inf_{x\in\mathbb{R}^{n}}\,\sup_{a\in A}\,\|x-a\|.$
@@
Definition A.1 (Trainable Super-Expressive Activation Function).
where $\sigma^{\star}:\mathbb{R}\to\mathbb{R}$ is given by
A.2 Proof Sketch
Lemma A.2.
@@ A.2 Proof Sketch
At each node of the tree, we will place an MLP which only locally approximates the target function on a little ball of suitable radius (implied by the tree valency $v$ and height $h$) of lemma A.2. I.e. by the storage space we would like to allocate to our neural pathway model. The next step of the proof relies on a mild extension of the quantitative universal approximation theorem in Shen et al. (2022a); Lu et al. (2021a) to the multivariate case, as well as an extension of the multivariate approximation result of Acciaio et al. (2023, Proposition 3.10) beyond the Hölder case.
Lemma A.3 (Universal Approximation Theorem).
Let $n,m\in\mathbb{N}_{+}$ with $n\geq 3$, ${\mathcal{K}}\subseteq\mathbb{R}^{n}$ be compact set $f:{\mathcal{K}}\to\mathbb{R}^{m}$ be uniformly continuous with strictly monotone continuous modulus of continuity $\omega$. Let $\sigma$ be an activation function as in Definitions A.1 or 3.1. For each $\varepsilon>0$, there exists an MLP $\hat{f}_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with trainable activation function $\sigma$ satisfying the uniform estimate
@@
Table 5: Complexity of the MLP $\hat{f}_{\theta}$ in Lemma A.3.
Activation $\sigma$ Super Expressive A.1 PReLU 3.1
Depth $(J)$ $15m$ $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\operatorname{rad}(K)2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
Width ($\max_{j}\,d_{j}$) $36n(2n+1)+m$ $16\max\{n,3\}+m$
Combining Lemmata A.2 and A.3 we obtain Theorem 3.4.
Lemma A.4.
$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
We have the following estimates:
  1. (i)
    Depth: Depth of each $\hat{f}_{\theta_{i}}$ is $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
  2. (ii)
    Width: Width of each $\hat{f}_{\theta_{i}}$ is $16\,\max\{n,3\}+m$
  3. (iii)
  4. (iv)
  5. (v)
where $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
We now present the technical version of Theorem 3.4. This result allows distributed neural computing using $\nu$-ary trees and allows for the approximation general uniformly continuous target functions.
Theorem A.5 (Technical Version of Theorem 3.4).
Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$ and a uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$. Fix an “approximation rate” $r\geq 0$, a “valency parameter” $v\in\mathbb{N}$ with $v\geq 2$, and an “approximation error” $\varepsilon>0$. Suppose that $\sigma$ satisfies Definition 3.1. Then, there is an $v$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ satisfying
$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\frac{\varepsilon^{-2r/n}}{2}\,\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}.$
Furthermore, for each $x\in{\mathcal{K}}$ and $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
@@
Table 6: Complexity of Feedforward Neural Network $\hat{f}_{\theta}$ and the $\nu$-ary tree in Theorem A.5. Here $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
@@
Leaves $\mathcal{O}\Big{(}v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\varepsilon^{2r/n}\,\operatorname{diam}({\mathcal{K}}))/\big{(}2\,\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)}\big{)}\big{\rceil},}\Big{)}$
Height $\big{\lceil}c\,\log(C)\big{(}1+\log(\epsilon^{2r/n}\operatorname{diam}({\mathcal{K}})/\big{(}2\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)})\big{)}\big{\rceil}$
@@
Appendix B Appendix: Proof of Main Theorem
Proof of Lemma A.2.
@@ Appendix B Appendix: Proof of Main Theorem
By Jung’s Theorem (see (Jung, 1901)), there exists $x_{0}\in\mathbb{R}^{n}$ such that the closed Euclidean ball $\overline{\operatorname{Ball}_{(\mathbb{R}^{n},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$ contains ${\mathcal{K}}$. Therefore, by Hölder’s inequality, we have that the $n$-dimensional hypercube $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$   [^3] For $x,y\in\mathbb{R}^{n}$ we denote by $[x,y]$ the hypercube defined by $\prod_{i=1}^{n}[x_{i},y_{i}]$.contains $\overline{B_{({\mathbb{R}^{n}},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$, where $\bar{1}=(1,\dots,1)\in\mathbb{R}^{n}$. Consequently, ${\mathcal{K}}\subseteq[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$. Let $\tilde{f}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}F|_{[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]}$, then $\tilde{f}\in C([x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}],\mathbb{R}^{m})$ is an $\omega$-uniformly continuous extension of $f$ to $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$.
@@
    If $\sigma$ satisfies Definition A.1 then, setting each $\alpha=0$ implies that $\sigma_{0}(x)=\sigma^{\star}(x)$, as defined in Eq. (10); thus
@@
If the activation function $\sigma$ is either in Definitions A.1 or 3.1, then it trivially implements the identity $I_{\mathbb{R}}$ on $\mathbb{R}$ by setting $\alpha=0$; i.e. $\sigma_{1}=I_{\mathbb{R}}$. Consequentially, for any $k\in\mathbb{N}_{+}$, if $I_{k}$ denotes the $k\times k$-identity matrix, then $I_{k}\sigma_{1}\bullet I_{k}\in\mathcal{NN}_{[d_{k}]}^{\sigma}$ with $P([d])=2k$, and $I_{k}\sigma_{1}\bullet I_{k}=1_{\mathbb{R}^{k}}$. Therefore, mutatis mutandis, $\mathcal{NN}_{[\cdot]}^{\sigma}$ satisfies the $c$-identity requirement with [^4] Formally, it satisfies what is the $1$-identity requirement, thus it satisfies the $c$-identity requirement for all integers $c\geq 2$. However, the authors of Cheridito et al. (2021) do not explicitly consider the extremal case where $c=1$. $c=2$, as defined in Cheridito et al. (2021, Definition 4). From there, mutatis mutandis, we may apply Cheridito et al. (2021, Proposition 5). Thus, there is a multi-index $[d]=(d_{0},\dots,d_{J})$ with $d_{0}=n$ and $d_{J}=m$, and a network $\hat{f}_{\theta}\in\mathcal{NN}_{[d]}^{\sigma}$ implementing $\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}e_{i}$, i.e.
@@
    If $\sigma$ satisfies Definition A.1 then, setting each $\alpha=0$
@@
Proof of Lemma A.4.
Consider the tree $\tilde{\mathcal{T}}$ given by Lemma A.2. For each leaf $v_{i}$ of $\tilde{\mathcal{T}}$, we apply Lemma A.3 to deduce the existence of an MLP $\hat{f}_{\theta_{i}}$ with explicit depth and width estimates given by that lemma, satisfying the uniform estimate
@@
We now prove our main technical version, namely Theorem A.5, which directly implies the special case recorded in Theorem 3.4.
Proof of Theorem A.5 (and this Theorem 3.4).
Applying Lemma A.4 with $\delta$ given as the solution of
@@
Setting ${\mathcal{K}}=B_{n}(0,1)$, $r=1/2$, $\omega(t)=Lt$, and thus $\omega^{-1}(t)=L^{-1}\,t^{1/\alpha}$, in Theorem A.5 yields the conclusion. Finally, by the computation in Example 1, we have that $C\leq 2^{n+1}$; thus, $\log(C)=(n+1)\log(2)\leq 2n$. Noting that $c=1/\log(2)=1$ completes the proof. ∎
Remark B.1.
[…]
B.1 Proofs of Results Concerning VC dimension
@@
We may now derive Proposition 3.6 by merging Lemma B.2 and one of the main results of Bartlett et al. (2019).
Proof of Proposition 3.6.

### v2

@@ 5 Overview of Derivation
We now overview the proof of our main result and its full technical formulation. These objectives require us to recall definitions from the analysis of metric spaces, which were not required in the statement of our main result but which are required when overviewing our proof.
5.1 Technical Definitions
@@
The metric ball in $(\mathcal{X},d)$ of radius $r>0$ at $x\in\mathcal{X}$ is denoted by $\operatorname{Ball}_{(\mathcal{X},d)}(x,r)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{z\in\mathcal{X}\colon d(x,z)<r\}$. A metric space $(\mathcal{X},d)$ is called doubling, if there is $C\in\mathbb{N}_{+}$ for which every metric ball in $(\mathcal{X},d)$ can be covered by at most $C$ metric balls of half its radius. The smallest such constant is called $(\mathcal{X},d)$’s doubling number, and is here denoted by $C_{(\mathcal{X},d)}$. Though this definition may seem abstract at first, Heinonen (2001, Theorem 12.1) provides an almost familiar characterization of all doubling metric spaces; indeed, $\mathcal{K}$ is a doubling metric space if and only if it can be identified via a suitable invertible map [^5] So called quasi-symmetric maps, see (Heinonen, 2001, page 78). with a subset of some Euclidean space. Every subset of $\mathbb{R}^{n}$, for any $n\in\mathbb{N}_{+}$, is a doubling metric space; see Robinson (2011, Chapter 9) for details.
@@
Fix a dimension $n\in\mathbb{N}_{+}$. The doubling number of any subset of Euclidean space is [^6] See Robinson (2011, Lemma 9.2) and the brief computations made in the proof of Robinson (2011, Lemma 9.4). at most $2^{n+1}$.
In what follows, all logarithms will be taken base $2$, unless explicitly stated otherwise, i.e. $\log_{v}$ is base $v$ for a given $v\in\mathbb{N}_{+}$ and $\log\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log_{2}$. As in Petrova & Wojtaszczyk (2023, page 762), the radius of a subset $A\subseteq\mathbb{R}^{n}$, denoted by $\operatorname{rad}(A)$, is defined by
$\operatorname{rad}(A)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\inf_{x\in\mathbb{R}^{n}}\,\sup_{a\in A}\,\|x-a\|.$ (9)
@@
5.2 Helping to Explain MoEs via Proof Sketch
Lemma 5.1 (Size of a Tree Whose Nodes Form a $\delta$-net of a Compact Subset of $\mathbb{R}^{n}$).
@@ 5.2 Helping to Explain MoEs via Proof Sketch
At each node of the tree, we will place an MLP which only locally approximates the target function on a little ball of suitable radius (implied by the tree valency $v$ and height $h$) of lemma 5.1. I.e. by the storage space we would like to allocate to our MoMLP model. The next step of the proof relies on a mild extension of the quantitative universal approximation theorem in Shen et al. (2022a); Lu et al. (2021a) to the multivariate case, as well as an extension of the multivariate approximation result of Acciaio et al. (2023, Proposition 3.10) beyond the Hölder case.
Lemma 5.2 (Vector-Valued Universal Approximation Theorem with Explicit Diameter Dependence).
Let $n,m\in\mathbb{N}_{+}$ with $n\geq 3$, ${\mathcal{K}}\subseteq\mathbb{R}^{n}$ be compact set of radius $\delta\geq 0$, $f:{\mathcal{K}}\to\mathbb{R}^{m}$ be uniformly continuous with strictly monotone continuous modulus of continuity $\omega$. Let $\sigma$ be an activation function as in Definitions 4.3 or 3.1. For each $\varepsilon>0$, there exists an MLP $\hat{f}_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with trainable activation function $\sigma$ satisfying the uniform estimate
@@
Table 3: Complexity of the MLP $\hat{f}_{\theta}$ in Lemma 5.2. See Table 5 in Appendix A for more detailed estimates.
Activation $\sigma$ Super Expressive 4.3 PReLU 3.1
Depth $(J)$ $\mathcal{O}(1)$ $\mathcal{O}\big{(}(\delta/\omega^{-1}(\varepsilon))^{n/2}\big{)}$
Width ($\max_{j}\,d_{j}$) $\mathcal{O}(1)$ $\mathcal{O}(1)$
Combining Lemmata 5.1 and 5.2 we obtain Theorem 4.1. We now present the technical version of Theorem 4.1. This result allows distributed neural computing using $\nu$-ary trees and allows for the approximation general uniformly continuous target functions.
Theorem 5.3 (Trade-Off: No. Expert vs. Expert Complexity - Technical Version of Theorem 4.1).
Suppose that $\sigma$ satisfies Definition 3.1. Let ${\mathcal{K}}$ be a compact subset of $\mathbb{R}^{n}$ whose doubling number is $C$. Fix an “number of experts-to-expert complexity trade-off parameter” $r\in\mathbb{R}$ and a “valency parameter” $\nu\in\mathbb{N}$ with $\nu\geq 2$.
For every uniformly continuous map $f:{\mathcal{K}}\to\mathbb{R}^{m}$ with modulus of continuity $\omega$ and each approximation error $\varepsilon>0$, $p\in\mathbb{N}_{+}$, there is an $\nu$-ary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ and a family of MLPs with (P)ReLU activation function $\{\hat{f}_{\theta_{i}}\}_{i=1}^{L}$ defined by $p$ parameters mapping $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ satisfying:
$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|<\frac{\varepsilon^{-2r/n}}{2}\,\omega^{-1}\biggl{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\biggr{)}$
and for each $x\in{\mathcal{K}}$ and $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
@@
Table 4: Complexity of Feedforward Neural Network $\hat{f}_{\theta}$ and the $\nu$-ary routing tree in Theorem 5.3. Here $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
Parameter Estimate
Depth $(J)$ $\mathcal{O}(\max\{1,\varepsilon^{-r}\})$
Width ($\max_{j}\,d_{j}$) $\mathcal{O}(1)$
No. Experts $\mathcal{O}\big{(}\max\{1,\varepsilon^{2r/n}/\omega^{-1}(\varepsilon)\}\big{)}$
Routing Complexity $\mathcal{O}\big{(}\max\{1,\log(\varepsilon^{2r/n}/\omega^{-1}(\varepsilon))\}\big{)}$
See Table 6 in Appendix A for more detailed estimates.
6 Conclusion
We presented approximation-theoretic and statistical foundations for MoEs by analysing the MoMLP model. We found that MoMLPs can indeed achieve arbitrary uniform approximation accuracy of continuous functions on compact subsets of Euclidean space while maintaining a feasible number of parameters in active GPU VRAM memory (Theorem 5.3). However, this naturally comes at the cost of requiring an exponential number of experts. We also obtain upper bounds on the VC dimension of the MoMLP model (Theorem 4.2), akin to the results of Bartlett et al. (2019) for ReLU MLPs, showing that deep-learning-based MoEs can indeed generalize.
Limitations and Future Work
This paper provides theoretical support, both in terms of approximation theory (approximation rates) and learning theory (VC dimension bounds), of MoE models. As such there is no clear limitation to our analysis. In future work, one can use the tools developed here to similarly study MoEs of other (deep) learning models.
Societal Impact
This paper presents work aimed at advancing the theoretical foundations of machine learning. Therefore, the majority of societal consequences are more positive than negative, as they acknowledge the necessity for principled machine learning and contribute to the development of understanding the capabilities of AI.
Acknowledgment and Funding
The authors thank James Lucas, Rafid Mahmood, and Gabriel Conant for their helpful feedback while preparing the manuscript.
AK acknowledges financial support from an NSERC Discovery Grant No. RGPIN-2023-04482 and their McMaster Startup Funds. AK also acknowledges that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute https://vectorinstitute.ai/partnerships/current-partners/. HSOB acknowledges financial support from an EV travel grant.
Appendix A Detailed Tables And Rates
Table 5: Complexity of the MLP $\hat{f}_{\theta}$ in Lemma 5.2.
Activation $\sigma$ Super Expressive 4.3 PReLU 3.1
[…]
Width ($\max_{j}\,d_{j}$) $36n(2n+1)+m$ $16\max\{n,3\}+m$
Table 6: Complexity of Feedforward Neural Network $\hat{f}_{\theta}$ and the $\nu$-ary tree in Theorem 5.3. Here $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
@@ Appendix A Detailed Tables And Rates
No. Experts (No. Leaves) $\mathcal{O}\Big{(}v^{\big{\lceil}c\,\log(C)\big{(}1+\log(\varepsilon^{2r/n}\,\operatorname{diam}({\mathcal{K}}))/\big{(}2\,\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)}\big{)}\big{\rceil},}\Big{)}$
Height (Routing Complexity) $\big{\lceil}c\,\log(C)\big{(}1+\log(\epsilon^{2r/n}\operatorname{diam}({\mathcal{K}})/\big{(}2\omega^{-1}\big{(}\frac{\varepsilon}{131\,(nm)^{1/2}}\big{)}\big{)})\big{)}\big{\rceil}$
@@
Appendix B Appendix: Proofs
B.1 Proof of Theorem 4.1
Proof of Lemma 5.1.
@@ Appendix B Appendix: Proofs
By Jung’s Theorem (see Jung (1901)), there exists $x_{0}\in\mathbb{R}^{n}$ such that the closed Euclidean ball $\overline{\operatorname{Ball}_{(\mathbb{R}^{n},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$ contains ${\mathcal{K}}$. Therefore, by Hölder’s inequality, we have that the $n$-dimensional hypercube $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$   [^7] For $x,y\in\mathbb{R}^{n}$ we denote by $[x,y]$ the hypercube defined by $\prod_{i=1}^{n}[x_{i},y_{i}]$.contains $\overline{B_{({\mathbb{R}^{n}},d_{n})}\left(x_{0},r_{\mathcal{K}}\right)}$, where $\bar{1}=(1,\dots,1)\in\mathbb{R}^{n}$. Consequently, ${\mathcal{K}}\subseteq[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$. Let $\tilde{f}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}F|_{[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]}$, then $\tilde{f}\in C([x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}],\mathbb{R}^{m})$ is an $\omega$-uniformly continuous extension of $f$ to $[x_{0}-r_{\mathcal{K}}\bar{1},x_{0}+r_{\mathcal{K}}\bar{1}]$.
@@
    If $\sigma$ satisfies Definition 4.3 then, setting each $\alpha=0$ implies that $\sigma_{0}(x)=\sigma^{\star}(x)$, as defined in Eq. (8); thus
@@
If the activation function $\sigma$ is either in Definitions 4.3 or 3.1, then it trivially implements the identity $I_{\mathbb{R}}$ on $\mathbb{R}$ by setting $\alpha=0$; i.e. $\sigma_{1}=I_{\mathbb{R}}$. Consequentially, for any $k\in\mathbb{N}_{+}$, if $I_{k}$ denotes the $k\times k$-identity matrix, then $I_{k}\sigma_{1}\bullet I_{k}\in\mathcal{NN}_{[d_{k}]}^{\sigma}$ with $P([d])=2k$, and $I_{k}\sigma_{1}\bullet I_{k}=1_{\mathbb{R}^{k}}$. Therefore, mutatis mutandis, $\mathcal{NN}_{[\cdot]}^{\sigma}$ satisfies the $c$-identity requirement with [^8] Formally, it satisfies what is the $1$-identity requirement, thus it satisfies the $c$-identity requirement for all integers $c\geq 2$. However, the authors of Cheridito et al. (2021) do not explicitly consider the extremal case where $c=1$. $c=2$, as defined in Cheridito et al. (2021, Definition 4). From there, mutatis mutandis, we may apply Cheridito et al. (2021, Proposition 5). Thus, there is a multi-index $[d]=(d_{0},\dots,d_{J})$ with $d_{0}=n$ and $d_{J}=m$, and a network $\hat{f}_{\theta}\in\mathcal{NN}_{[d]}^{\sigma}$ implementing $\sum_{i=1}^{m}\hat{f}_{\theta^{(i)}}e_{i}$, i.e.
@@
    If $\sigma$ satisfies Definition 4.3 then, setting each $\alpha=0$
@@
Lemma B.1 (Trade-Off: No. Expert vs. Expert Complexity).
$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
We have the following estimates:
  1. (i)
    Depth. Depth of each $\hat{f}_{\theta_{i}}$ is $m\left(19+2n+11\Big{\lceil}\Big{(}\frac{\delta 2^{3/2}n^{1/2}}{(n+1)^{1/2}\omega^{-1}\big{(}\varepsilon/(131\sqrt{n\,m})\big{)}}\Big{)}^{n/2}\Big{\rceil}\right)$
  2. (ii)
    Width. Width of each $\hat{f}_{\theta_{i}}$ is $16\,\max\{n,3\}+m$
  3. (iii)
  4. (iv)
  5. (v)
where $c\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\log(v)^{-1}$.
Proof of Lemma B.1.
Consider the tree $\tilde{\mathcal{T}}$ given by Lemma 5.1. For each leaf $v_{i}$ of $\tilde{\mathcal{T}}$, we apply Lemma 5.2 to deduce the existence of an MLP $\hat{f}_{\theta_{i}}$ with explicit depth and width estimates given by that lemma, satisfying the uniform estimate
@@
We now prove our main technical version, namely Theorem 5.3, which directly implies the special case recorded in Theorem 4.1.
Proof of Theorem 5.3 (and this Theorem 4.1).
Applying Lemma B.1 with $\delta$ given as the solution of
@@
Setting ${\mathcal{K}}=B_{n}(0,1)$, $r=1/2$, $\omega(t)=Lt$, and thus $\omega^{-1}(t)=L^{-1}\,t^{1/\alpha}$, in Theorem 5.3 yields the conclusion. Finally, by the computation in Example 1, we have that $C\leq 2^{n+1}$; thus, $\log(C)=(n+1)\log(2)\leq 2n$. Noting that $c=1/\log(2)=1$ completes the proof. ∎
B.2 Proofs of Theorem 4.2
@@
We may now derive Theorem 4.2 by merging Lemma B.2 and one of the main results of Bartlett et al. (2019).
Proof of Theorem 4.2.